OpenStudy (anonymous):

so 10 to power of zero is zero

OpenStudy (mathteacher1729):

Anything to the power zero is one. EXCEPT zero to the zero, that is NOT DEFINED.

OpenStudy (anonymous):

how can ten to (0) power equal 1

OpenStudy (anonymous):

It just does.

OpenStudy (mathteacher1729):

It is defined that anything to the zero power is one. We can derive it from the quotient rules for powers: \[\huge \frac{x^a}{x^b}=x^{a-b}\] So let x = 10, and let a = b (so that a - b = 0). \[\huge 1=\frac{10^4}{10^4}=10^{4-4}=10^0=1\] :)

OpenStudy (phi):